OpenStudy (anonymous):

Crystal has 70 ribbons. Fourteen of them are red. What percent of her ribbons are red?

OpenStudy (anonymous):

14/70 = ?

OpenStudy (anonymous):

0.2

OpenStudy (anonymous):

Yep which translates to 20%

OpenStudy (anonymous):

:O, you are a guenus

OpenStudy (triciaal):

percent means per 100 set up the ratio 14/70 = x/100 where x is the unknown percentage

OpenStudy (triciaal):

yes 20%